See Exercise 9.15 in the Chapter 9 Programming Exercise from the Book section for the description of this exercise.​

Sample Run

Enter the real part of the first complex number: 3.5

Enter the imaginary part of the first complex number: 6.5

Enter the real part of the first complex number: -3.5

Enter the imaginary part of the first complex number: 1

(3.5 + 6.5i) + (-3.5 + 1i) = (0.0 + 7.5i)

(3.5 + 6.5i) - (-3.5 + 1i) = (7.0 + 5.5i)

(3.5 + 6.5i) * (-3.5 + 1i) = (-18.75 - 19.25i)

(3.5 + 6.5i) / (-3.5 + 1i) = (-0.43396226415 - 1.98113207547i)

|(3.5 + 6.5i)| = 4.47213595499958.

In Python.
Answer #1

Code:

import math
#taking inputs for two complex numbers
r1=input("Enter the real part of the first complex number: ")
i1=input("Enter the imaginary part of the first complex number: ")
r2=input("Enter the real part of the second complex number: ")
i2=input("Enter the imaginary part of the second complex number: ")

#addition of two complex numbers
def add(r1,i1,r2,i2):
return r1+r2,i1+i2

#substraction of two complex numbers
def sub(r1,i1,r2,i2):
return r1-r2,i1-i2

#multiplication of two complex numbers
def mul(r1,i1,r2,i2):
return (r1*r2-i1*i2),(r1*i2+i1*r2)

#division of two complex numbers
def div(r1,i1,r2,i2):
r=(r1*r2+i1*i2)/(r2*r2+i2*i2)
i=(i1*r2-r1*i2)/(r2*r2+i2*i2)
return r,i

#absolute value of complex number
def abslt(r,i):
return math.sqrt(r*r+i*i)

#printing result
def prnt(r1,i1,r2,i2,r,i,c):
print("(%.2f + %.2fi) %c (%.2f + %.2fi) = (%.2f + %.2fi)")%(r1,i1,c,r2,i2,r,i)

r,i=add(r1,i1,r2,i2)
prnt(r1,i1,r2,i2,r,i,'+')

r,i=sub(r1,i1,r2,i2)
prnt(r1,i1,r2,i2,r,i,'-')

r,i=mul(r1,i1,r2,i2)
prnt(r1,i1,r2,i2,r,i,'*')

r,i=div(r1,i1,r2,i2)
prnt(r1,i1,r2,i2,r,i,'/')

val=abslt(r1,i1)
print("|(%.2f + %.2fi)| = %f")%(r1,i1,val)


OUTPUT:

Enter the real part of the first complex number: 3.5 Enter the imaginary part of the first complex number: 5.4 Enter the real